OpenStudy (babyslapmafro):

Please help me find a value for m so that y=x^m is a solution of the given differential equation. (x^2)y''-y=0

OpenStudy (loser66):

since y =x\(^m\) is solution, we have y'= m x\(^{m-1}\) y"= m (m-1)x \(^{m-2}\) x\(^2\)y"= m(m-1)(m-2)x\(^m\) Yours becomes m(m-1)(m-2)x\(^m\)-x\(^m\)=0 factor x \(^m\) out, solve for m you have value of such that y = x\(^m\) is solution.

OpenStudy (loser66):

*value of m such that...

OpenStudy (anonymous):

@Babyslapmafro, got it?

OpenStudy (babyslapmafro):

why does (m-2) become a coefficient?

OpenStudy (babyslapmafro):

@SithsAndGiggles @loser66

OpenStudy (anonymous):

Minor typo, I think: \[\begin{align*} y&=x^m\\ y'&=mx^{m-1}\\ y''&=m(m-1)x^{m-2}&\Rightarrow~~x^2y''=m(m-1)x^m \end{align*}\] Substituting into the equation you have \[m(m-1)x^m-x^m=0\\ x^m\left(m^2-m-1\right)=0\]